Find the values of t that bound the middle 0.8 of the distribution for df = 30. (Give your answers correct to two decimal places.)
------to-----
Answer #1

solution :

Given that,

Degrees of freedom = 30

Using standard normal table,

P( -t < t < t) = 0.8

= P(t < t) - P(t <-t ) = 0.8

= 2P(t < t) - 1 = 0.8

= 2P(t < t) = 1 + 0.8

= P(t < t) = 1.8 / 2

= P(t < t) = 0.9

= P(t < 1.310) = 0.9

= t ± 1.31